OpenStudy (anonymous):

Help please. I have a final tomorrow and am struggling with chapter 3. Use the given coniditions to write an equation for each line in point-slope form and slope intercept form. Slope = -2 passing through (-1,4)

OpenStudy (anonymous):

ok then m= slope b= y intercept point slope form looks like this \[y_2-y_1=m(x_2-x_1)\] or \[y-y_1=m(x-x_1)\] slope intercept form looks like \[y=mx+b\]

OpenStudy (anonymous):

got those formulas but having trouble using them

OpenStudy (anonymous):

if you're given the slope of the line and a point, substitute those values into the point slope formula then convert that into slope intercept if you need to

OpenStudy (precal):